The sides of a parallelogram have lengths 8 inches and 16 inches and one of the angles of the parallelogram has a measure of 45 degrees. Find the area of the parallelogram?
Answer by checkley75(3666)

THE ENDS OF THE PARALLELOGRAM FORM A RIGHT ISOSOLESE TRIANGLE WITH SIDES X, X & 8. 8 BEING THE HYPOTENUSE. WE NEED TO FIND THE OTHER SIDE FOR THE HEIGHT OF THE PARALLELOGRAM THUS:
X^2+X^2=8^2
2X^2=64
X^2=64/2
X^2=32
X=5.65685 THIS IS THE OTHER TWO SIDES & THE HEIGHT OF THE PARALLELOGRAM.
NOW WE MULTIPLY THE BASE TIMES THE HEIGHT.
16*5.65685=90.51 IN^2